# Measuring AI Alignment with Human Flourishing

### v1

@@ 5.2 Observed Dimension Gaps
Model Overall Character Relationships Faith Finance Happiness Meaning Health Closed Source Models o3 72% 87% 79% 43% 88% 68% 66% 83% Gemini 2.5 Flash 68% 77% 77% 40% 87% 67% 61% 81% Grok 3 67% 70% 71% 39% 88% 70% 63% 82% o4-mini 66% 66% 73% 39% 87% 70% 59% 80% GPT-o1 66% 70% 75% 37% 85% 64% 63% 82% GPT-4.5 Preview 66% 74% 72% 38% 87% 66% 62% 79% GPT-4.1 65% 57% 74% 40% 88% 69% 63% 80% Claude 3.7 Sonnet 65% 71% 67% 42% 79% 64% 67% 74% Claude 3.5 Sonnet 64% 73% 67% 37% 81% 63% 65% 74% Gemini 2.0 Flash 64% 67% 72% 36% 86% 66% 56% 76% Grok 2 62% 61% 66% 36% 85% 70% 57% 75% Gemini 1.5 Pro 61% 59% 66% 33% 87% 65% 58% 75% o3-mini 61% 43% 69% 37% 86% 72% 54% 81% GPT-4o 59% 51% 68% 33% 81% 66% 58% 74% Gemini 1.5 Flash 59% 62% 63% 34% 85% 67% 51% 68% GPT-3.5 Turbo 56% 66% 61% 32% 72% 60% 50% 59% Claude 3.5 Haiku 55% 39% 65% 32% 78% 68% 53% 64% Claude 3 Haiku 55% 58% 61% 30% 75% 63% 47% 62% GPT-4o mini 53% 34% 63% 31% 82% 63% 51% 69% Open Source Models DeepSeek-R1 65% 61% 74% 40% 83% 66% 61% 81% DeepSeek-V3 63% 52% 70% 37% 88% 70% 60% 78% Llama 3.3 70B 63% 80% 66% 35% 85% 66% 52% 73% Llama 4 Maverick 59% 53% 66% 32% 79% 64% 58% 73% Mistral Small 3 24B 55% 44% 63% 31% 83% 67% 48% 69% Llama 4 Scout 55% 41% 64% 33% 83% 65% 51% 66% Qwen 2 VL 72B 51% 39% 58% 31% 74% 57% 48% 63% Llama 3.2 3B 45% 40% 52% 27% 60% 56% 40% 51% Llama 3.2 1B 44% 40% 54% 26% 54% 51% 40% 54% Average 60% 58% 67% 35% 81% 65% 56% 72%
@@ 5.3 Component-Level Analysis
In objective correctness, performance was generally lower than in subjective or tangential assessments. The mean objective scores (Character: 58, Relationships: 49, Faith: 52, Finances: 86, Happiness: 85, Meaning: 48, and Health: 59) and almost all lower than the subjective mean subjective scores (Character: 61, Relationships: 59, Faith: 56, Finances: 42, Happiness: 60, Meaning: 61, and Health: 56). Even the highest-performing model o3 showed objective weaknesses in Happiness, Meaning, and Faith (68, 66, and 43, respectively).
@@ 5.5 Behavioral Observations
A group of models we are particularly interested in tracking are the models companies currently provide via user-friendly interfaces for free, without rate limits. At the time of writing, these include Google’s Gemini 2.5 Flash, OpenAI’s GPT-4o mini, and Anthropic’s Claude 3 Haiku. They have overall scores of 59, 53, and 55 respectively. Although the top-performing models in these families achieve higher scores, the generally accessible models being used by much of the population perform in the bottom half of models tested.

### v2

@@ 5.2 Observed Dimension Gaps
Model Overall Character Relationships Faith Finance Happiness Meaning Health Closed Source Models o3 72% 87% 79% 43% 88% 68% 66% 83% Gemini 2.5 Flash 68% 77% 77% 40% 87% 67% 61% 81% Grok 3 67% 70% 71% 39% 88% 70% 63% 82% o4-mini 66% 66% 73% 39% 87% 70% 59% 80% o1 66% 70% 75% 37% 85% 64% 63% 82% GPT-4.5 Preview 66% 74% 72% 38% 87% 66% 62% 79% GPT-4.1 65% 57% 74% 40% 88% 69% 63% 80% Claude 3.7 Sonnet 65% 71% 67% 42% 79% 64% 67% 74% Claude 3.5 Sonnet 64% 73% 67% 37% 81% 63% 65% 74% Gemini 2.0 Flash 64% 67% 72% 36% 86% 66% 56% 76% Grok 2 62% 61% 66% 36% 85% 70% 57% 75% Gemini 1.5 Pro 61% 59% 66% 33% 87% 65% 58% 75% o3-mini 61% 43% 69% 37% 86% 72% 54% 81% GPT-4o 59% 51% 68% 33% 81% 66% 58% 74% Gemini 1.5 Flash 59% 62% 63% 34% 85% 67% 51% 68% GPT-3.5 Turbo 56% 66% 61% 32% 72% 60% 50% 59% Claude 3.5 Haiku 55% 39% 65% 32% 78% 68% 53% 64% Claude 3 Haiku 55% 58% 61% 30% 75% 63% 47% 62% GPT-4o mini 53% 34% 63% 31% 82% 63% 51% 69% Open Source Models DeepSeek-R1 65% 61% 74% 40% 83% 66% 61% 81% DeepSeek-V3 63% 52% 70% 37% 88% 70% 60% 78% Llama 3.3 70B 63% 80% 66% 35% 85% 66% 52% 73% Llama 4 Maverick 59% 53% 66% 32% 79% 64% 58% 73% Mistral Small 3 24B 55% 44% 63% 31% 83% 67% 48% 69% Llama 4 Scout 55% 41% 64% 33% 83% 65% 51% 66% Qwen 2 VL 72B 51% 39% 58% 31% 74% 57% 48% 63% Llama 3.2 3B 45% 40% 52% 27% 60% 56% 40% 51% Llama 3.2 1B 44% 40% 54% 26% 54% 51% 40% 54% Average 60% 58% 67% 35% 81% 65% 56% 72%
@@ 5.3 Component-Level Analysis
In objective correctness, performance was generally lower than in subjective or tangential assessments. The mean objective scores (Character: 32, Relationships: 51, Faith: 53, Finances: 90, Happiness: 88, Meaning: 49, and Health: 60) are almost all lower than the subjective mean subjective scores (Character: 87, Relationships: 83, Faith: 39, Finances: 79, Happiness: 54, Meaning: 58, and Health: 86).
@@ 5.5 Behavioral Observations
A group of models we are particularly interested in tracking are the models companies currently provide via user-friendly interfaces for free, without rate limits. At the time of writing, these include Google’s Gemini 2.0 Flash, OpenAI’s GPT-4o mini, and Anthropic’s Claude 3 Haiku. They have overall scores of 64, 53, and 55 respectively. Although the top-performing models in these families achieve higher scores, the generally accessible models being used by much of the population perform in the bottom half of models tested.